# Descent-Guided Policy Gradient for Scalable Cooperative Multi-Agent Learning

### v1

@@ Abstract
Scaling cooperative multi-agent reinforcement learning (MARL) is fundamentally limited by cross-agent noise: when agents share a common reward, the actions of all $N$ agents jointly determine each agent’s learning signal, so cross-agent noise grows with NN. In the policy gradient setting, per-agent gradient estimate variance scales as $\Theta(N)$, yielding sample complexity $\mathcal{O}(N/\epsilon)$. We observe that many domains—cloud computing, transportation, power systems—have differentiable analytical models that prescribe efficient system states. In this work, we propose Descent-Guided Policy Gradient (DG-PG), a framework that constructs noise-free per-agent guidance gradients from these analytical models, decoupling each agent’s gradient from the actions of all others. We prove that DG-PG reduces gradient variance from $\Theta(N)$ to $\mathcal{O}(1)$, preserves the equilibria of the cooperative game, and achieves agent-independent sample complexity $\mathcal{O}(1/\epsilon)$. On a heterogeneous cloud scheduling task with up to 200 agents, DG-PG converges within 10 episodes at every tested scale—from N=5N{=}5 to N=200N{=}200—directly confirming the predicted scale-invariant complexity, while MAPPO and IPPO fail to converge under identical architectures.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) has demonstrated remarkable success across a wide range of domains, from game playing (Yu et al., 2022) to cloud resource scheduling (Mao et al., 2019) and transportation management (El-Tantawy et al., 2013). In many of these applications, agents share a common objective and must coordinate their behavior—a setting known as cooperative MARL. However, scaling cooperative MARL remains a fundamental challenge: performance degrades sharply as the number of agents grows. The root of this difficulty is multi-agent credit assignment: when all agents share a single reward signal, it becomes increasingly hard to determine how much each agent’s action contributed to the observed outcome, and this ambiguity worsens with every additional agent.
Concretely, any optimization signal in cooperative MARL—a policy gradient, a Q-value, or an advantage estimate—must be derived from a shared return that aggregates the actions of all $N$ agents, introducing cross-agent noise that grows linearly with NN. Kuba et al. (2021) formalize this in the policy gradient setting, showing that the variance of each agent’s gradient estimate scales as $\Theta(N)$—implying sample complexity $\mathcal{O}(N/\epsilon)$ and a signal-to-noise ratio that decays as $1/\sqrt{N}$. The central bottleneck in scaling cooperative MARL is thus the cross-agent noise inherent to learning from shared rewards.
Existing approaches address cross-agent noise through two broad strategies. The first is credit decomposition: extracting per-agent contributions from the shared return via counterfactual baselines (Foerster et al., 2018), value factorization (Rashid et al., 2020; Wang et al., 2021a), or marginal contribution estimation (Wolpert and Tumer, 2001). These methods can reduce but not eliminate cross-agent noise, as they remain dependent on the shared return. The second strategy uses pre-designed controllers as base policies (Levine and Koltun, 2013; Johannink et al., 2019), but constrains the learned policy to remain near the base controller, potentially biasing it away from the true optimum. Neither approach provides per-agent learning signals that are both noise-free and unbiased.
An alternative source of noise-free per-agent signals does exist: many cooperative problems—cloud scheduling, traffic routing, power dispatch—have differentiable analytical models developed over decades in operations research and control theory. These models typically rely on simplifying assumptions and global system information, and thus cannot be directly translated into executable policies for individual agents. However, their differentiable structure can provide each agent with a noise-free guidance gradient that depends only on the agent’s own influence on the system state.
@@
  • Framework. We augment the cooperative objective with a guidance term derived from analytical models. The resulting gradient decomposes into a standard policy gradient term and a per-agent guidance gradient computed directly from the analytical model; because this term depends only on each agent’s own influence on the system state, it is free of cross-agent noise. The framework requires no architectural changes to existing actor-critic methods.
  • Theoretical guarantees. We prove three properties under mild structural assumptions: (i) the augmented objective preserves the stationary points of the original cooperative game (Nash invariance, Theorem 4.1); (ii) the per-agent gradient estimate variance reduces from $\Theta(N)$ to $\mathcal{O}(1)$, independent of the number of agents (Theorem 4.2); and (iii) the sample complexity to reach an $\epsilon$-optimal solution is $\mathcal{O}(1/\epsilon)$, also independent of $N$ (Theorem 4.3).
@@ 2.1 Cooperative Objective and Policy Gradient
$\nabla_{\theta^{i}}J(\boldsymbol{\theta})=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}|o_{t}^{i})\,G_{t}\right],$ (2)
where $G_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}$ is the sampled return from time $t$.
@@ 2.2 The Variance Explosion Problem
Existing variance reduction techniques, such as state-value baselines $V(s)$ or counterfactual baselines like COMA (Foerster et al., 2018), primarily target the self-exploration term $\sigma_{\text{self}}^{2}$. While effective in smaller settings, their ability to mitigate the $\Theta(N)$ cross-agent noise term diminishes in large-scale systems without incurring significant computational overhead. Similarly, mean-field methods (Yang et al., 2018) offer scalability but require agents to be exchangeable and homogeneous—assumptions that do not hold in domains such as cloud scheduling, power grids, and traffic networks, where agents have heterogeneous capacities and structured coupling constraints.
@@ 3 Method: Descent-Guided Policy Gradient
In this section, we present DG-PG, a framework that addresses the variance explosion problem by exploiting the analytical models available in many cooperative domains. We first introduce a reference state—an efficient system state prescribed by analytical models—and the conditions under which it yields a valid guidance gradient (Section 3.1). We then construct an augmented objective whose gradient decomposes into a standard policy gradient term and a noise-free, per-agent guidance gradient derived from the reference (Section 3.2). Under mild assumptions, this formulation preserves the stationary points of the original cooperative objective while substantially reducing gradient estimate variance.
3.1 Leveraging Analytical Priors: System State and Reference
In many cooperative domains, analytical models from operations research and control theory—such as fluid-limit approximations for queueing systems, equilibrium analysis for traffic networks, or convex optimization for resource allocation—provide principled characterizations of desirable system behavior. While not executable under real-world stochasticity, these models offer something valuable: a reference that describes what efficient system operation looks like under given conditions. Our key observation is that such a reference, if properly defined, can supply each agent with a clear, directed signal for policy improvement—one that is far more informative than the noisy scalar reward $r_{t}$ aggregated over all agents.
As noted above, domain-specific models can prescribe efficient system states—but these states must be expressed in a concrete representational space. We define a system state $\boldsymbol{x}_{t}\in\mathbb{R}^{d}$ as the vector of system-level quantities jointly determined by agents’ actions—for example, per-server resource utilization, per-link flow, or queue lengths. The reference state $\tilde{\boldsymbol{x}}_{t}\in\mathbb{R}^{d}$ is the efficient state that the analytical model yields under current conditions, computed independently of the learned policies. A key structural property of $\boldsymbol{x}_{t}$ is that each agent’s action typically affects only specific dimensions of this vector, preserving the locality of the problem—a property we exploit in the gradient computation (Section 3.2).
@@ 3.1 Leveraging Analytical Priors: System State and Reference
When both conditions hold, the reference provides a stable, directionally correct target: it does not shift as policies update (Assumption 3.1), and approaching it is expected to improve the cooperative objective (Assumption 3.2). This means that any mechanism that steers the system state toward $\tilde{\boldsymbol{x}}_{t}$ is aligned with the original optimization goal—a property we exploit in the next subsection to construct a low-variance gradient estimator. Both assumptions are broadly satisfied in structured cooperative domains and can be approximately verified in practice; we demonstrate this concretely for the cloud scheduling setting in Appendix C.
@@
Importantly, $\tilde{\boldsymbol{x}}_{t}$ need not be reachable by the agents—it may arise from idealized fluid-limit or steady-state models. DG-PG uses $\tilde{\boldsymbol{x}}_{t}$ only to define a descent direction, not as a target to be attained exactly.
3.2 Descent-Guided Formulation
@@
Given a reference $\tilde{\boldsymbol{x}}_{t}$ satisfying Assumptions 3.13.2, we know that steering the system toward $\tilde{\boldsymbol{x}}_{t}$ is directionally aligned with improving $J$. The challenge is to incorporate this reference into the learning process in a way that (i) does not alter the optimal policies of the original cooperative game, (ii) reduces the $\Theta(N)$ gradient variance identified in Section 2.2, and (iii) provides each agent with a localized signal about its own contribution. We achieve all three by augmenting the objective with a differentiable guidance functional and exploiting its structure in the gradient computation.
Augmented Objective. We define a guidance functional $\mathcal{G}$ that measures deviation from the reference:
$\mathcal{G}(\boldsymbol{\pi})\triangleq\mathbb{E}_{\tau\sim\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}d(\boldsymbol{x}_{t},\tilde{\boldsymbol{x}}_{t})\right],$ (6)
with $d(\boldsymbol{x}_{t},\tilde{\boldsymbol{x}}_{t})=\frac{1}{2}\|\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t}\|^{2}$. By Assumption 3.2, minimizing $\mathcal{G}$ is directionally consistent with maximizing the cooperative return $J$, so we combine them into an augmented objective:
$\max_{\boldsymbol{\pi}}J_{\alpha}(\boldsymbol{\pi})\triangleq(1-\alpha)J(\boldsymbol{\pi})-\alpha\mathcal{G}(\boldsymbol{\pi}),$ (7)
where $\alpha\in(0,1)$ controls the balance. By linearity of expectation, this is equivalent to optimizing a shaped per-step reward
$r_{t}^{\text{aug}}=(1-\alpha)\,r_{t}-\alpha\cdot d(\boldsymbol{x}_{t},\tilde{\boldsymbol{x}}_{t})$
with any standard policy gradient method. However, naively using $r_{t}^{\text{aug}}$ as a shaped reward would still estimate the gradient via $\nabla\log\pi\cdot r_{t}^{\text{aug}}$: since $r_{t}^{\text{aug}}$ depends on all agents’ actions, the estimator variance remains $\Theta(N)$.
Instead, DG-PG exploits the differentiable structure of $\mathcal{G}$ directly: as we show next, its gradient decomposes into per-agent terms that can be computed analytically, without relying on sampled returns. Moreover, the augmentation preserves the original stationary points because the guidance gradient vanishes at optimality (Theorem 4.1).
Gradient Decomposition. The gradient of $J_{\alpha}$ separates into two terms:
$\nabla_{\boldsymbol{\theta}}J_{\alpha}=(1-\alpha)\nabla_{\boldsymbol{\theta}}J-\alpha\nabla_{\boldsymbol{\theta}}\mathcal{G}.$ (8)
The first term is the standard policy gradient, whose coefficient $G_{t}$ aggregates the stochastic effects of all $N$ agents and must be estimated via Monte Carlo rollouts, yielding $\Theta(N)$ variance. The second term, however, admits an entirely different estimation strategy. Because $d(\boldsymbol{x}_{t},\tilde{\boldsymbol{x}}_{t})$ is a known, differentiable function of the system state with gradient $\nabla_{\boldsymbol{x}}d=\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t}$, we can differentiate through $\boldsymbol{x}_{t}$ directly via the chain rule, without relying on sampled returns. Applying this for agent $i$:
@@
$\nabla_{\theta^{i}}\mathcal{G}$ $=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}\left\langle\nabla_{\boldsymbol{x}}d(\boldsymbol{x}_{t},\tilde{\boldsymbol{x}}_{t}),\frac{\partial\boldsymbol{x}_{t}}{\partial a^{i}_{t}}\right\rangle\nabla_{\theta^{i}}\log\pi^{i}(a^{i}_{t}|o^{i}_{t})\right]$
$=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\boldsymbol{z}_{t}^{i}\rangle\nabla_{\theta^{i}}\log\pi^{i}(a^{i}_{t}|o^{i}_{t})\right],$ (9)
@@
where $\boldsymbol{z}_{t}^{i}=\frac{\partial\boldsymbol{x}_{t}}{\partial a^{i}_{t}}$ is agent $i$’s local influence vector—the Jacobian of the system state with respect to its action. The concrete form of $\boldsymbol{z}_{t}^{i}$ depends on the domain; we derive it for cloud scheduling in Appendix D.
Intuitively, the guidance coefficient $\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\boldsymbol{z}_{t}^{i}\rangle$ measures the system deviation projected onto the dimensions that agent $i$ actually influences. Unlike $G_{t}$, this coefficient does not depend on other agents’ actions: $\boldsymbol{x}_{t}$ is observed, $\tilde{\boldsymbol{x}}_{t}$ is exogenous, and $\boldsymbol{z}_{t}^{i}$ captures only agent $i$’s own influence on the system state. As $\alpha$ increases, the estimator shifts weight from the stochastic policy gradient term toward this deterministic guidance term, reducing variance. We formalize this in Section 4.
Gradient Estimator. Combining both terms, the complete DG-PG gradient estimator for agent $i$ is:
$\hat{g}_{\text{DG}}^{i}$ $=(1-\alpha)\,\hat{g}_{J}^{i}$
@@
$\quad-\alpha\sum_{t=0}^{T}\gamma^{t}\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\,\boldsymbol{z}^{i}_{t}\rangle\,\nabla_{\theta^{i}}\log\pi^{i}(a^{i}_{t}|o^{i}_{t}).$ (10)

### v2

@@ Abstract
Scaling cooperative multi-agent reinforcement learning (MARL) is fundamentally limited by cross-agent noise. When agents share a common reward, the actions of all $N$ agents jointly determine each agent’s learning signal, so cross-agent noise grows with NN. In the policy gradient setting, per-agent gradient estimate variance scales as $\Theta(N)$, yielding sample complexity $\mathcal{O}(N/\epsilon)$. We observe that many domains, including cloud computing, transportation, and power systems, have differentiable analytical models that prescribe efficient system states. In this work, we propose Descent-Guided Policy Gradient (DG-PG), a framework that utilizes these analytical models to provide each agent with a noise-free gradient signal, decoupling each agent’s gradient from the actions of all others. We prove that DG-PG reduces gradient variance from $\Theta(N)$ to $\mathcal{O}(1)$, preserves the equilibria of the cooperative game, and achieves agent-independent sample complexity $\mathcal{O}(1/\epsilon)$. On a heterogeneous cloud scheduling task with up to 200 agents, DG-PG converges within 10 episodes at every tested scale, from N=5N{=}5 to N=200N{=}200, directly confirming the predicted scale-invariant complexity, while MAPPO and IPPO fail to converge under identical architectures.
@@ 1 Introduction
Multi-agent reinforcement learning (MARL) has demonstrated remarkable success across a wide range of domains, from game playing (Yu et al., 2022) to cloud resource scheduling (Mao et al., 2019) and transportation management (El-Tantawy et al., 2013). In many of these applications, agents share a common objective and must coordinate their behavior, a setting known as cooperative MARL. However, scaling cooperative MARL remains a fundamental challenge, as performance degrades sharply with the number of agents. The root of this difficulty is multi-agent credit assignment: when all agents share a single reward signal, it becomes increasingly hard to determine how much each agent’s action contributed to the observed outcome, and this ambiguity worsens with every additional agent.
Concretely, any optimization signal in cooperative MARL, whether a policy gradient, a Q-value, or an advantage estimate, must be derived from a shared return that aggregates the actions of all $N$ agents, introducing cross-agent noise that grows linearly with NN. Kuba et al. (2021) formalize this in the policy gradient setting, showing that the variance of each agent’s gradient estimate scales as $\Theta(N)$, implying sample complexity $\mathcal{O}(N/\epsilon)$ and a signal-to-noise ratio that decays as $1/\sqrt{N}$. The central bottleneck in scaling cooperative MARL is thus the cross-agent noise inherent to learning from shared rewards.
Existing approaches attempt to mitigate cross-agent noise through two broad strategies. The first is credit decomposition, extracting per-agent contributions from the shared return via counterfactual baselines (Foerster et al., 2018), value factorization (Rashid et al., 2020; Wang et al., 2021), or marginal contribution estimation (Wolpert and Tumer, 2001). These methods reduce but do not eliminate cross-agent noise, as the learning signal remains derived from the shared return. The second strategy incorporates pre-designed controllers as base policies (Levine and Koltun, 2013; Johannink et al., 2019), providing a noise-free action component. However, this constrains the learned policy near the controller, which may prevent it from reaching the true optimum. In both cases, cross-agent noise is either reduced but not eliminated, or avoided at the cost of constraining the solution space.
We observe that many cooperative systems already possess an underexploited source of noise-free per-agent signals. For example, in domains such as cloud scheduling, traffic routing, and power dispatch, differentiable analytical models have been developed over decades in operations research and control theory. These models typically rely on simplifying assumptions and global system information, and thus cannot be directly translated into executable policies for individual agents. However, their differentiable structure can provide each agent with a noise-free gradient signal that depends only on the agent’s own influence on the system state.
@@
  • Framework. We augment the policy gradient with the gradient of a differentiable guidance term derived from analytical models. Crucially, the differentiable structure of these models allows this gradient to isolate each agent’s individual contribution to the system state, yielding a noise-free signal computable without any other agent’s action. The framework requires no architectural changes to existing actor-critic methods.
  • Theoretical guarantees. We prove three properties under mild structural assumptions: (i) the augmented gradient does not alter the Nash equilibria of the original cooperative game (Theorem 4.1); (ii) the per-agent gradient estimate variance reduces from $\Theta(N)$ to $\mathcal{O}(1)$, independent of the number of agents (Theorem 4.2); and (iii) the sample complexity to reach an $\epsilon$-optimal solution is $\mathcal{O}(1/\epsilon)$, also independent of $N$ (Theorem 4.3).
@@ 2.1 Cooperative Objective and Policy Gradient
$\nabla_{\theta^{i}}J(\boldsymbol{\theta})=\mathbb{E}_{\boldsymbol{\pi}}\left[\sum_{t=0}^{\infty}\gamma^{t}\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a_{t}^{i}|o_{t}^{i})\,R_{t}\right],$ (2)
where $R_{t}=\sum_{k=0}^{\infty}\gamma^{k}r_{t+k}$ is the sampled return from time $t$.
@@ 2.2 The Variance Explosion Problem
Existing variance reduction techniques, such as state-value baselines $V(s)$ or counterfactual baselines like COMA (Foerster et al., 2018), primarily target the self-exploration term $\sigma_{\text{self}}^{2}$. While effective in smaller settings, their ability to mitigate the $\Theta(N)$ cross-agent noise term diminishes in large-scale systems without incurring significant computational overhead. Similarly, mean-field methods (Yang et al., 2018) offer scalability but require agents to be exchangeable and homogeneous, assumptions that do not hold in domains such as cloud scheduling, power grids, and traffic networks, where agents have heterogeneous capacities and structured coupling constraints.
@@ 3 Method: Descent-Guided Policy Gradient
In this section, we present DG-PG, a framework that addresses the variance explosion problem by exploiting the analytical models available in many cooperative domains. We first define the reference state as an efficient system state prescribed by analytical models and identify the conditions under which it yields a valid guidance term (Section 3.1). We then augment the policy gradient with a noise-free, per-agent signal derived from the reference (Section 3.2). Under mild assumptions, DG-PG does not alter the Nash equilibria of the original cooperative game while achieving agent-independent variance.
3.1 Utilizing Analytical Priors: System State and Reference
In many cooperative domains, analytical models from operations research and control theory, such as fluid-limit approximations for queueing systems, equilibrium analysis for traffic networks, or convex optimization for resource allocation, provide principled characterizations of desirable system behavior. While not executable under real-world stochasticity, these models offer something valuable, namely a reference that describes what efficient system operation looks like under given conditions. Our key observation is that such a reference, if properly defined, can supply each agent with a clear, directed signal for policy improvement. This signal is far more informative than the noisy scalar reward $r_{t}$ aggregated over all agents.
As noted above, domain-specific models can prescribe efficient system states, but these states must be expressed in a concrete representational space. We define a system state $\boldsymbol{x}_{t}\in\mathbb{R}^{m}$ as the vector of system-level quantities jointly determined by agents’ actions (e.g., per-server resource utilization, per-link flow, or queue lengths), where $m$ is the number of monitored quantities. The reference state $\tilde{\boldsymbol{x}}_{t}\in\mathbb{R}^{m}$ is the efficient state that the analytical model yields under current conditions, computed independently of the learned policies. A key structural property of $\boldsymbol{x}_{t}$ is that each agent’s action typically affects only specific dimensions of this vector, so that each agent’s influence on the system state is localized. We exploit this property in the gradient computation (Section 3.2).
@@ 3.1 Utilizing Analytical Priors: System State and Reference
When both conditions hold, the reference provides a stable, directionally correct target that does not shift as policies update (Assumption 3.1), and approaching it is expected to improve the cooperative objective (Assumption 3.2). This means that any mechanism that steers the system state toward $\tilde{\boldsymbol{x}}_{t}$ is aligned with the original optimization goal, a property we exploit in the next subsection to construct a low-variance gradient estimator. Both assumptions are broadly satisfied in structured cooperative domains and can be approximately verified in practice. We verify both assumptions concretely for the cloud scheduling setting in Appendix C.
@@
Importantly, $\tilde{\boldsymbol{x}}_{t}$ need not be reachable by the agents, because it may arise from idealized fluid-limit or steady-state models. DG-PG uses $\tilde{\boldsymbol{x}}_{t}$ only to define a descent direction, not as a target to be attained exactly.
3.2 Descent-Guided Policy Gradient
@@
Given a reference $\tilde{\boldsymbol{x}}_{t}$ satisfying Assumptions 3.13.2, steering the system toward $\tilde{\boldsymbol{x}}_{t}$ at each step is directionally aligned with improving $J$. The challenge is to incorporate this information into the learning process in a way that (i) does not alter the original cooperative objective, and (ii) provides each agent with a gradient signal free of cross-agent noise. Our DG-PG achieves both by exploiting the differentiable structure of these models at the gradient level, providing each agent with its own noise-free update direction.
Concretely, we construct the DG-PG framework in three steps. We first define a deviation measure that quantifies the gap between the current system state and the reference. We then differentiate this measure through each agent’s action to obtain a per-agent descent direction. Finally, we use these directions to guide the standard policy gradient, thus eliminating cross-agent noise.
Deviation Measure. We first define a per-step deviation function to quantify how far the system state deviates from the reference:
$d(\boldsymbol{x}_{t},\tilde{\boldsymbol{x}}_{t})=\frac{1}{2}\|\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t}\|^{2}.$ (6)
By Assumption 3.2, reducing $d$ at each step is directionally consistent with improving $J$. However, simply incorporating $d$ into the reward would yield a scalar signal that cannot distinguish each agent’s individual effect. Since $d$ is differentiable with respect to $\boldsymbol{x}_{t}$ and $\boldsymbol{x}_{t}$ depends differentiably on each $a_{t}^{i}$, the gradient $\partial d/\partial a_{t}^{i}$ can be computed analytically. Thus, the gradient of $d$ can serve as a guidance signal for each agent, aligned with improving $J$.
DG-PG Gradient. We augment the standard policy gradient with the gradient of $d$, yielding the DG-PG gradient:
$\nabla_{\theta^{i}}J_{\alpha}\triangleq(1-\alpha)\,\nabla_{\theta^{i}}J-\alpha\,\nabla_{\theta^{i}}\mathcal{G},$ (7)
where $\alpha\in(0,1)$ is the guidance weight, and $\mathcal{G}\triangleq\mathbb{E}_{s\sim\nu^{\boldsymbol{\pi}}}\!\left[d(\boldsymbol{x}_{t},\tilde{\boldsymbol{x}}_{t})\right]$ is the expected deviation under the state visitation distribution $\nu^{\boldsymbol{\pi}}$ induced by the joint policy. The DG-PG gradient $\nabla_{\theta^{i}}J_{\alpha}$ thus retains the optimization direction of $J$ while incorporating a guidance signal from the reference model. We prove that this augmentation preserves the Nash equilibria of the original game (Theorem 4.1).
To evaluate $\nabla_{\theta^{i}}\mathcal{G}$, we apply the chain rule through $\boldsymbol{x}_{t}$ to isolate each agent’s contribution to the deviation:
@@
$\frac{\partial d}{\partial a_{t}^{i}}=\left\langle\nabla_{\boldsymbol{x}}d,\,\frac{\partial\boldsymbol{x}_{t}}{\partial a^{i}_{t}}\right\rangle=\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\,\boldsymbol{z}_{t}^{i}\rangle,$ (8)
@@
where $\boldsymbol{z}_{t}^{i}\triangleq\frac{\partial\boldsymbol{x}_{t}}{\partial a^{i}_{t}}$ is agent $i$’s local influence vector, the partial derivative of the system state with respect to its action. This coefficient measures the deviation projected onto the dimensions that agent $i$ influences. It depends only on the current state $\boldsymbol{x}_{t}$, the exogenous reference $\tilde{\boldsymbol{x}}_{t}$, and the agent-local derivative $\boldsymbol{z}_{t}^{i}$, so it carries no cross-agent noise and can be computed analytically. The concrete form of $\boldsymbol{z}_{t}^{i}$ depends on the domain, and we derive it for cloud scheduling in Appendix D.
Substituting Equation 8 into the policy gradient theorem yields the gradient of the guidance term:
@@
$\nabla_{\theta^{i}}\mathcal{G}=\mathbb{E}_{s\sim\nu^{\boldsymbol{\pi}},\,\boldsymbol{a}\sim\boldsymbol{\pi}}\left[\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\,\boldsymbol{z}_{t}^{i}\rangle\,\nabla_{\theta^{i}}\log\pi^{i}(a^{i}_{t}|o^{i}_{t})\right].$ (9)
Gradient Estimator. In practice, we estimate the DG-PG gradient $\nabla_{\theta^{i}}J_{\alpha}$ via the score function method. At each step $t$ of a sampled trajectory, the gradient estimate for agent $i$ is:
$\hat{g}_{\text{DG},t}^{i}=(1-\alpha)\,\hat{g}_{J,t}^{i}-\alpha\,\hat{g}_{\mathcal{G},t}^{i},$ (10)
where
$\hat{g}_{J,t}^{i}$ $=R_{t}\,\nabla_{\theta^{i}}\log\pi^{i}(a^{i}_{t}|o^{i}_{t}),$ (11)
$\hat{g}_{\mathcal{G},t}^{i}$ $=\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\,\boldsymbol{z}^{i}_{t}\rangle\,\nabla_{\theta^{i}}\log\pi^{i}(a^{i}_{t}|o^{i}_{t}).$ (12)
Each agent then updates its policy parameters via gradient ascent:
$\theta^{i}_{k+1}=\theta^{i}_{k}+\eta\,\hat{g}_{\text{DG},t}^{i},$ (13)
where $\eta>0$ is the learning rate.
Variance Properties. The two components of $\hat{g}_{\text{DG},t}^{i}$ have fundamentally different variance properties. The standard policy gradient sample $\hat{g}_{J,t}^{i}$ uses the return $R_{t}$, which aggregates all $N$ agents’ stochastic actions, yielding $\Theta(N)$ variance. The guidance sample $\hat{g}_{\mathcal{G},t}^{i}$ uses an analytically computed coefficient $\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\boldsymbol{z}^{i}_{t}\rangle$, contributing $\mathcal{O}(1)$ variance independent of $N$. We formalize how the choice of $\alpha$ controls this variance trade-off in Section 4.